OpenStudy (testing123guest):

Which ordered pairs in the form (x, y) are solutions to the equation 7x – 5y = 28? Choose all answers that are correct. A. (−6, −14) B. (−1, −7) C. (4, 10) D. (7, 9)

OpenStudy (testing123guest):

how

OpenStudy (anonymous):

B

OpenStudy (anonymous):

just substitute each pair into the equation and check if it is correct.

OpenStudy (anonymous):

^

OpenStudy (mathstudent55):

Just plug in each choice into the equation and see if it makes a true statement. If it does, it satisfies the equation.

OpenStudy (anonymous):

the left side as to equal 28, put in all the numbers and find which one equals 28

OpenStudy (testing123guest):

i cant if i get it wrong the first time that is it for that question

OpenStudy (anonymous):

The B one works: 7(-1) - 5(-7) = 28 -7 + 35 = 28 28 = 28

OpenStudy (mathstudent55):

Here is the worked out problem for choice A. 7x – 5y = 28 A. (-6. -14) 7(-6) - 5(-14) = 28 -42 + 70 = 28 28 = 28 Since 28 = 28 is true, that means (-6, -14) is a solution, so choice A works. Now follow the same procedure for each of the other choices.

OpenStudy (anonymous):

C and D don't work c. 7(4) - 5(10)=28 28 - 50 = 28 -22 = 28 This is not true. D. 7(7) - 5(9)=28 49 - 45 = 28 4=28 This is also not true. So only A and B work.

OpenStudy (anonymous):

d because i said so